OpenStudy (anonymous):

Is the number square root of 2x is equvilent to the square root of 2x

OpenStudy (unklerhaukus):

\[\sqrt{2}x\neq\sqrt{2x}\]

OpenStudy (anonymous):

is it true or false ?

OpenStudy (anonymous):

true or false ?

OpenStudy (unklerhaukus):

try x=2 \[\sqrt{2}\times2=2\sqrt2\]\[\neq\]\[\sqrt{2\times2}=\sqrt4=2\]